OpenStudy (anonymous):

WILL GIVE SHINY MEDAL 3x less than or equal to 21 A) x less than or equal to 7 B) x greater than or equal to 7 C) x less than or equal to 18 D) x greater than or equal to 18

sammixboo (sammixboo):

\(\rm \color{#20bd23}{3x \le 21}\) All you need to to is divide both sides of the equation by 3 \(\rm \color{#20bd23}{3x\div 3 \le 21\div 3}\) \(\rm \color{#20bd23}{x \le 21\div 3}\)

sammixboo (sammixboo):

Now what is \(\rm \color{#20bd23}{21\div 3}\)?

OpenStudy (anonymous):

7

sammixboo (sammixboo):

Correct, so \(\rm \color{#20bd23}{x \le 7}\)

sammixboo (sammixboo):

Or x is less than or equal to 7

OpenStudy (anonymous):

thx!